OpenStudy (anonymous):

\[\sqrt[3]{t-78}\] what is the domain in interval notation please

OpenStudy (tkhunny):

There are no restrictions on odd roots.

OpenStudy (anonymous):

[78, infinity)

OpenStudy (anonymous):

that doesnt work for me for some odd reason

OpenStudy (anonymous):

my software dosent accept [78,inf). I have tried

OpenStudy (anonymous):

possible notation mistake?

OpenStudy (anonymous):

one sec

OpenStudy (tkhunny):

You're not paying attention. There are no restriction on odd roots. None. Try \((-\infty,\infty)\)

OpenStudy (anonymous):

yea my bad, tkhunny is right

OpenStudy (anonymous):

yea tkhunny right

OpenStudy (anonymous):

thanks

OpenStudy (tkhunny):

:-) It was bound to happen, sometime!

OpenStudy (anonymous):

tkhunny, can you help me with about four more questions please

OpenStudy (tkhunny):

Only if you promise to pay attention! :-)

OpenStudy (anonymous):

ok i will

OpenStudy (anonymous):

my first question is... Find the center and radius of the circle whose equation is x^2−1x+y^2−7y−14=0.

OpenStudy (tkhunny):

This requires "Complete the Square". Have you done it?

OpenStudy (anonymous):

i have tried but am really clueless as to how

OpenStudy (tkhunny):

Very bad answer. We'll pick it up on the other thread.